y=-1 divided by 5-4
Answer by checkley77(12844)

y=-1 divided by 5-4
Do you mean:
y=(-1/5)-4
y=-1/5-20/5
y=-21/5 ans.
OR:
y=-1/(5-4)
y=-1/1
y=-1 ans.